Circular No. 8304

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2003 VB_{12}

A very-slow-moving body reported by M. E. Brown, C. A. Trujillo, and D. L. Rabinowitz, on CCD images obtained with the Palomar 1.2-m Schmidt telescope on 2003 Nov. 14, is apparently the second-largest known transneptunian object (after Pluto), with an absolute magnitude H = 1.7. Subsequent observations have been found on CCD images and sky-survey photographs back to 2001 by Brown et al., leading to orbital elements (T = 2076, q = 76 AU, Q ≈ 1000 AU, i = 11^{o}.9; cf. MPEC 2004-E45).

COMET C/2001 Q4 (NEAT) A. Lecacheux, N. Biver, J. Crovisier, and D. Bockelée-Morvan, Obser-

vatoire de Paris, on behalf of the Odin Team, write: “The H_{2}O 110-101 line at 556.936 GHz was observed in comet C/2001 Q4 on Mar. 6.6 UT with high spectral resolution (80 m/s). The line integrated intensity (main beam brightness temperature) was 2.7 ± 0.1 K km/s over the 2-hr observation. This corresponds to a preliminary water-production rate of 1.3 × 10^{29} molecules/s, taking into account opacity effects.”

Visual total-magnitude estimates by A. Pearce, Nedlands, W.Australia (20×80 binoculars): Mar. 9.55 UT, 7.5; 10.52, 7.4; 11.52, 7.4; 13.50, 7.3; 15.52, 7.1.

SUPERNOVAE 2004an, 2004ao, AND 2004at Further to IAUC 8303, T. Matheson, P. Challis, and R. Kirshner re-

port that a spectrum of SN 2004an (cf. IAUC 8299), obtained by P. Berlind on Mar. 14.45 UT, shows it to be a type-II supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 8979 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 8800 km/s. A spectrum of SN 2004ao (cf. IAUC 8299), obtained on Mar. 14.53, shows it to be a type-Ib supernova, similar to SN 1998dt at 8 days past maximum (see Matheson et al. 2001, A.J. 121, 1648). A spectrum of SN 2004at (cf. IAUC 8302), obtained on Mar. 15.53, shows it to be a type-II supernova.

IRAS 05436–0007

Corrigendum. On IAUC 8301, line 9, for Paδ read Paβ

2004 March 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green